Mod 12.3b: Note: The following is not a journal entry nor is it a full trial balance, it is just a random selection of entries to given accounts. Your job is to identify from the list of accounts below, which accounts would be classified as "Equity" accounts (i.e. accounts that directly impact overall equity). Once you have identified all Equity accounts, compute how the debits or credits in such accounts would impact a company's total equity. For example, if it would increase by 4 then just enter the number 4. If it would decrease by 4, enter the number as -4. Account name Debit Credit Unearned Sales Revenue 37 Bonds payable 21 Sales Revenue 51 Common Stock 17 Salaries payable 9 Dividends 4 Cash 14 The answer that I put was 13 but I got it wrong can you help me please?

accounting software
Answer #1

Answer:

Accounts directly impact Equity account Impact(increase/decrease)
Common stock 17
Sales revenue 51
Dividends -4
Net impact (17+51)-4 =68-4 = 64

The over all equity will increase by 64.

Explanation:

Equity = capital + retained earnings

1. Common stock is a capital account. It will increase equity.

2.Retained earnings = (Revenue - expenses) - dividends. Sales revenue will cause an increase in reatined earnings and dividends will decrease the retained earnings. That's how they affect equity account.

3. other accounts are of other categories. They can't influence equity account

Unearned sales revenue Current liability
Bonds payable Current liability(assumed to of short term nature)
Salaries payable Current liability
Cash Current asset